Consider the following acquisition data regarding Wellington Industries and Orators Telecom Inc.: Wellington Industries is considering an acquisition of Orators Telecom Inc. Wellington Industries estimates that acquiring Orators will result in incremental value for the fim. The analysts involved in the deal have collected the following information from the projected financial statements of the target company Data Collected (in millions of dollars) Year 1 EBIT Interest expense Debt Total net operating cap 4.0 34.1 119.5 Year 2 6.0 4.4 40.3 121.8 Year 3 7.5 4.8 ital 124.1 Orators is a publicly traded company, and its market-determined pre-merger beta is 1.20. You also have the following information about the company and the projected statements. .Orators currently has a $18.00 million market value of equity and $11.70 million in debt. * The risk-free rate is 5% with a 7.10% market risk premium, and the Capital Asset Pricing Model produces a pre-merger required rate of return on equity rSL of 13.52%. Oratorss cost of debt is 7.00% at a tax rate of 40%. * The projections assume that the company will have a post-horizon growth rate of 5.00%. .Current total net operating capital is $116.0, and the sum of existing debtand debt required to maintain a constant capital structure at the time of acquisition is $31 million The fim has no nonoperating assets, such as marketable securities. With the given information, use the free cash flow to equity (FCFE) approach to calculate the following values involved in the merger analysis Value FCFE horizon value Value of FCFE The estimated value of Orators Oratorss shareholders wil s operations after the merger IS than the market value of Oratorss equity. This means that the wealth of if it merges with Wellington rather than remaining as a stand-alone corporation.
Answer #1

Calculations

1.Horizon value of Unlevered cash flows = FCF3(1+g) / reu - g

where,

FCF3 is free cash flow for 3rd year, g is growth rate that is 5& and reu is unlevered Cost of equity which is 11.12%

reu = rf + (rm-rf)Bu

reu = 5+7.10*0.86

=11.12%

First Calculate Bu,

BU = BL/1+(1-T)*D/E

BU = 1.20/ 1+(.60)*11.70/18

BU =0.86

.

So, firstly we need to calculate FCF

Year 1

Year 2

Year 3

EBIT

5

6

7.5

Tax

2

2.4

3

NOPAT (EBIT-TAX)

3

3.6

4.5

Less: Net New Investment in Operating Capital

3.5

2.3

2.3

FCF

-0.05

1.3

2.2

As Horizon value of Unlevered cash flows = FCF3(1+g) / reu - g , we will insert the figures,

Horizon value of Unlevered cash flows = 2.2(1+.05)/0.1112-0.05)

= $37.75 million

2) Unlevered Value of Operations = FCF1/(1+reu) +FCF2/ (1+r eu)2 +FCF3 +UHV3/ (1+reu)3

=-0.05/1.1112 + 1.3/1.11122+ (2.2+37.75)/1.11123

= $30.12 million

3. Horizon Value of Tax Shield = ITS3(1+g)/reu -g

So, firstly we need to calculate ITS

  

Interest Expense

4

4.4

4.8

Interest Tax Saving (ITS) = Tax Rate*Interest Saving

1.6

1.76

1.92

As Horizon Value of Tax Shield = ITS3(1+g)/reu -g, we will insert the figures -

  Horizon Value of Tax Shield = 1.92 (1.05) /0.1112-0.05

= 32.94 million

4. Value of tax shield = ITS1/(1+reu) +ITS2/ (1+r eu)2 +ITS3 +HVITS3/ (1+reu)3

  =1.6/1.1112+1.76*1.11122+(1.92+32.94)/1.11123

  =0.94+ 0.92+ 15.057

=$28.27 million

5. Value of operation = Unlevered Value of Operations + Value of tax shield

= 30.12+28.27

= $ 58.39 million

Calculation of Total Value of Equity

As the firm has no non operating asset , the Value of operation is equal to Value of firm

VF= Ve+ Value of debt

Value of equity (Ve) =58.39-11.70

=$ 46.69 million

The Total Value of Orator's equity is $46.69 million

Fill ups

a)more

b) increase